OpenStudy (anonymous):

completwe the factoring if needed and solve the polynomial ineqaulity using a sign chart (2x+1)(x-2)(3x-4)≤0

OpenStudy (ash2326):

2x+1)(x-2)(3x-4)<=0 it can be written as (x+1/2)(x-2)(x-4/3)<=0 we have on the number line

OpenStudy (anonymous):

Your the best !

OpenStudy (ash2326):

wait i'm solving

OpenStudy (ash2326):

there are four ranges of interest a)x <=-1/2 b)-1/2<=x<=4/3 c) 4/3<=x<=2 d) x>=2 let's check the inequality for each range (x+1/2)(x-4/3)(x-2) a) when x<=-1/2 let's say x=-1 then (-1+1/2)(-1-4/3)(-1-2) -21/6 this is less than equal to 0 , so this is a part of our solution b) -1/2<=x<=4/3 let x = 1 then (1+1/2)(1-4/3)(1-2) 1/2 is not less than or equal to zero so this is not of our interest c) 4/3<=x<=2 let x be 3/2 (3/2+1/2)(3/2-4/3)(3/2-2) -1/6 <=0 this is a part of our solution d) when x>=2 if you evaluate it's positive son not of our interest so our solution is x<=-1/2 and 4/3<=x<=2